how do i find the equation of the parabola, y=ax^2+bx+c, that passes through the following three points : (-2,40),(1,7),(3,15)?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

how do i find the equation of the parabola, y=ax^2+bx+c, that passes through the following three points : (-2,40),(1,7),(3,15)?
----
Substitute those value for x and y.
The solve the tree equations for a,b,c.
==================
ax^2 + bx + c = y
a(4) + b(-2)+c = 40
a(1) + b(1) +c = 7
a(9) + b(3) +c = 15
----------------------------
I used a matrix method to solve the 3 equations and got:
a = 3
b = -8
c = 12
----
Equation:
y = 3x^2-8x+12
========================
Cheers,
Stan H.
=======================

Answer by Alan3354(69443)

Another method, using determinants:
----------
Parabola
x y x^2 Coeff
Point 1 -2 40 4 1
Point 2 1 7 1 1
Point 3 3 15 9 1
--------------
I have an Excel sheet that does these.
--> y = 3x^2 - 8x + 12
---------------------------
That was no help at all, except for confirming the solution.